OpenStudy (anonymous):

6x + 2y = 12 Can someone help me identify the slope and y-intercept?

OpenStudy (johnweldon1993):

Well we can make this into slope intercept form to make it easier y = mx + b 'm' is slope 'b' is y-intercept you have 6x + 2y = 12 Lets solve for 'y' subtract 6x from both sides 2y = -6x + 12 now divide everything by 2 to isolate 'y' 2y = -6x + 12 ------------- 2 2 2 y = -3x + 6 compare this to y = mx + b What is your slope...and what is your y-intercept...?

OpenStudy (anonymous):

my slope would be -3 and my y intercept would be 6? Is that right @johnweldon1993

OpenStudy (johnweldon1993):

That would be correct!

OpenStudy (anonymous):

Yay Thank you!!!!!!!!!!!!!!!!!!!!!u rock!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (johnweldon1993):

No problem :) and thank you! :)